In this question you should state the parameters of any distributions you use. An orchard sells kiwis of two varieties, Type A and Type B. The masses, in grams, of Type A and Type B kiwis follow the distributions N(110,62)N(110,6^2) and N(85,82)N(85,8^2) respectively. It is assumed that these two distributions are independent. Kiwis are selected randomly and packed into bags. Type A kiwis are packed in bags of 30, while Type B kiwis are packed in bags of 60. The selling prices of Type A and Type B kiwis are $25 per kilogram and $15 per kilogram respectively. The orchard also produces genetically modified Type G kiwis, which are packed in bags of 50. The masses of Type G kiwis are identically distributed with a mean of 90 grams and a standard deviation of 5 grams. Each Assortment Box contains one bag of Type A kiwis and one bag of Type G kiwis. A box is fit for sale if the total mass of the kiwis in the box is at least 7.85 kilograms.

  1. (a)

    In a randomly chosen bag of Type A kiwis, find the probability that the mean mass of kiwis is between 109 grams and 115 grams.

  2. (b)

    Find the probability that the average selling price of a kiwi from a bag of Type A kiwis is more than the average selling price of a kiwi from a bag of Type B kiwis by at most $1.50.

  3. (c)

    Ten Assortment Boxes are selected at random. Find the probability that none of the boxes are fit for sale.

  4. (d)

    State, in context, the approximation(s) and assumption(s) needed for the distribution(s) used in part (c).

Work through it with AI →

It won't just hand you the answer — it asks where you're stuck.How the checker works

Worked solution

10(a)

1. State the distribution of the sample mean mass of Type A kiwis.

Let XˉA\bar{X}_A be the sample mean mass of 30 Type A kiwis.
Since the mass of a Type A kiwi follows N(110,62)N(110, 6^2),
XˉA∼N(110,6230)\bar{X}_A \sim N\left(110, \frac{6^2}{30}\right)
XˉA∼N(110,1.2)\bar{X}_A \sim N(110, 1.2)

2. Calculate the probability that the sample mean is between 109 and 115 grams.

Using a graphing calculator,
P(109<XˉA<115)≈0.81934P(109 < \bar{X}_A < 115) \approx 0.81934
≈0.819 (to 3 s.f.)\approx 0.819 \text{ (to 3 s.f.)}

Answer

0.819

10(b)

1. Define the random variables for the average selling price of a kiwi from a bag of Type A and Type B kiwis.

Let Aˉ\bar{A} and Bˉ\bar{B} be the sample mean masses of kiwis in a bag of Type A (n=30n=30) and Type B (n=60n=60) respectively.
Let SAS_A and SBS_B be the average selling price of a kiwi from a bag of Type A and Type B kiwis.

The selling prices per gram are $251000=$0.025\frac{\char36{}25}{1000} = \char36{}0.025 for Type A and $151000=$0.015\frac{\char36{}15}{1000} = \char36{}0.015 for Type B.
Therefore, SA=0.025AˉS_A = 0.025\bar{A} and SB=0.015BˉS_B = 0.015\bar{B}.

2. Find the expectation and variance for the average selling price of a kiwi from a bag of Type A.

E(SA)=0.025×110=2.75\mathrm{E}(S_A) = 0.025 \times 110 = 2.75
Var(SA)=0.0252Var(Aˉ)=0.000625(6230)=0.00075\mathrm{Var}(S_A) = 0.025^2 \mathrm{Var}(\bar{A}) = 0.000625 \left(\frac{6^2}{30}\right) = 0.00075

3. Find the expectation and variance for the average selling price of a kiwi from a bag of Type B.

E(SB)=0.015×85=1.275\mathrm{E}(S_B) = 0.015 \times 85 = 1.275
Var(SB)=0.0152Var(Bˉ)=0.000225(8260)=0.00024\mathrm{Var}(S_B) = 0.015^2 \mathrm{Var}(\bar{B}) = 0.000225 \left(\frac{8^2}{60}\right) = 0.00024

4. Define the distribution for the difference in average selling prices and calculate the required probability.

Let D=SA−SBD = S_A - S_B.
E(D)=2.75−1.275=1.475\mathrm{E}(D) = 2.75 - 1.275 = 1.475
Var(D)=Var(SA)+Var(SB)=0.00075+0.00024=0.00099\mathrm{Var}(D) = \mathrm{Var}(S_A) + \mathrm{Var}(S_B) = 0.00075 + 0.00024 = 0.00099

D∼N(1.475,0.00099)D \sim \mathrm{N}(1.475, 0.00099)

We require the probability that SAS_A is more than SBS_B by at most $1.50\char36{}1.50, which means 0<D⩽1.500 < D \leqslant 1.50.
P(0<D⩽1.50)≈0.78656≈0.787\mathrm{P}(0 < D \leqslant 1.50) \approx 0.78656 \approx 0.787 (to 3 s.f.)

Answer

0.7870.787

10(c)

1. Find the distribution, mean, and variance of the total mass of an Assortment Box

Let MM be the total mass of the kiwis in an Assortment Box, consisting of 30 Type A kiwis and 50 Type G kiwis.

M=∑i=130Ai+∑j=150GjM = \sum_{i=1}^{30} A_i + \sum_{j=1}^{50} G_j

E(M)=30(110)+50(90)=7800\text{E}(M) = 30(110) + 50(90) = 7800
Var(M)=30(62)+50(52)=1080+1250=2330\text{Var}(M) = 30(6^2) + 50(5^2) = 1080 + 1250 = 2330
Since the sample size of Type G kiwis (n=50n=50) is large, by the Central Limit Theorem, their total mass follows an approximate normal distribution. Thus, the total mass MM is approximately normally distributed:
M∼N(7800,2330)M \sim \text{N}(7800, 2330)

2. Calculate the probability that a single Assortment Box is fit for sale

A box is fit for sale if its total mass is at least 7.85 kg=7850 g7.85 \text{ kg} = 7850 \text{ g}.
Using the normal distribution M∼N(7800,2330)M \sim \text{N}(7800, 2330):
P(M≥7850)=0.149954P(M \ge 7850) = 0.149954

3. Model the number of fit boxes with a binomial distribution and calculate the final probability

Let YY be the number of Assortment Boxes fit for sale out of 10 randomly selected boxes.
Y∼B(10,0.149954)Y \sim \text{B}(10, 0.149954)
We require the probability that none of the boxes are fit for sale (Y=0Y = 0):
P(Y=0)=(1−0.149954)10=0.19698≈0.197 (to 3 s.f.)P(Y = 0) = (1 - 0.149954)^{10} = 0.19698 \approx 0.197 \text{ (to 3 s.f.)}

Answer

0.197

10(d)

1. State the approximation needed for the distribution of the mass of Type G kiwis.

Approximation: The distribution of the masses of Type G kiwis is not given. However, since the sample size of 5050 Type G kiwis in a bag is sufficiently large (n≥30n \ge 30), we use the Central Limit Theorem to approximate the total mass of the 5050 Type G kiwis as being normally distributed.

2. State the assumption needed regarding the relationship between the masses of the kiwis.

Assumption: The question states that the masses of Type G kiwis are 'identically distributed' but does not explicitly state they are independent. We must assume that the mass of any Type G kiwi is independent of the mass of other Type G kiwis so that we can sum their variances and apply the Central Limit Theorem.
(Another valid assumption: The total mass of the bag of Type A kiwis is independent of the total mass of the bag of Type G kiwis in the same Assortment Box).

Answer

Approximation: Since the sample size of 5050 is large (≥30\ge 30), by the Central Limit Theorem, the total mass of the 5050 Type G kiwis is approximately normally distributed.

Assumption: The mass of any Type G kiwi is independent of the mass of any other Type G kiwi. (Alternatively: The masses of the Type A kiwis and Type G kiwis are independent of one another.)

AI analysis — may contain mistakes.

Worked this one out on paper? Snap your steps and Gigle's AI math checker will circle the step where it went wrong.

More questions on this topic

  1. 1Fishing Company A has two types of fishing vessels, the "Standard" vessel and the "Large" vessel. The amount of fish cau…
  2. 2Each day a train travels from Town A to Town B and it is scheduled to arrive at Town B at 0800 hrs. During the first yea…
  3. 3Korean dramas, or Kdramas, have been immensely popular among students, captivating their attention during leisure hours.…
  4. 4In this question you should clearly state the values of the appropriate parameters of any distributions used. The diamet…
  5. 5During training, the time in seconds for a soldier to dismantle a certain type of equipment is a normally distributed co…